Please help!!
x=1+6 sqrt x-9
Thank you!

Answer by fractalier(6550)

Okay, from x = 1 + 6sqrt(x-9), move the one over first and get
x - 1 = 6sqrt(x-9)
Now square both sides
x^2 - 2x + 1 = 36(x - 9)
Now expand, combine, factor, solve and check...
x^2 - 2x + 1 = 36x - 324
x^2 - 38x + 325 = 0
(x - 13)(x - 25) = 0
x = 13 or x = 25
Both check out, so we have two solutions...